Find a value for k so that 9m^2-kn^2 will have the factors 3m+7n and 3m-7n.
Thanks for your help.

Answer by venugopalramana(3286)

Find a value for k so that 9m^2-kn^2 will have the factors 3m+7n and 3m-7n.
(3m+7n)*(3m-7n)= (3M)^2- (7N)^2 = 9M^2-49N^2
COMPARING WITH 9m^2-kn^2...WE GET K=49